OpenStudy (anonymous):

solve the equation by expressing each side as a power of the same base and then equating exponents. 4^(3x+5)=1/256

OpenStudy (campbell_st):

well I think you need a base of 4... \[\frac{1}{256} = 4^{-4}\] so the equation can be written as \[(4)^{3x + 5} = 4^{-4}\] so now you just need to solve the powers for x 3x + 5 = -4 hope it helps

OpenStudy (anonymous):

yes that will give the answer I am looking for so to convert 1/256 I would use a negative power